What is the molarity of a solution that contains 0.350 g of barium fluoride, BaF2, in 0.50 L of solution?
1
Expert's answer
2013-01-17T10:37:47-0500
Molarity or C = mol per L

So you need to find amount of BaF2:

n = m/Mw ( Mw = 175)

n = 0.350/175 = 0.002 mol

So it is 0.002 mol per 0.50 L and concentration is two
much more becose it is some amount per one L

C = 0.004 M